OpenStudy (anonymous):

can anyone help me find the general solution to the 2nd order differential equation y'' − 5y' + 6y = 3

OpenStudy (anonymous):

the auxillary eqn for complementry fuction is m^2-5m+6=0 (m-3)(m-2)=0 m=2,3

OpenStudy (anonymous):

\[y _{c}=C _{1}e ^{2}x+C _{2}e ^{3}x\]

OpenStudy (anonymous):

\[y[x]=\frac{1}{2}+e^{2 x} C[1]+e^{3 x} C[2] \]

OpenStudy (anonymous):

This is a an undetermined coefficient type problem in case you need to go read your book and review.

OpenStudy (anonymous):

mathgirl, Are you talking to moi?

OpenStudy (anonymous):

no darlin, to raetp

OpenStudy (anonymous):

robtobey i get the auxilery function and how that works i just dont understand how the =3 part comes into play

OpenStudy (anonymous):

A commented Mathamatica solution is attached.